Fourth Attempt - Please show step-by-step calculations and what line items are used for the following ratio calculation so it is easy to understand what numbers have been used:

Note:Accounts Receivable Numbers used in 2nd and 3rd attempt that don't work are: Barco $37,400 and Kyan $57,400; 3rd attempt - $33600 - Barco and Kyan - $55,800. I have the remaining numbers that are correct.

Summary information from the financial statements of two companies competing in the same industry follows. Barco Company Kyan Company Barco Company Kyan Company Data from the current year-end balance sheets Data from the current year’s income statement Assets Sales $ 770,000 $ 880,200 Cash $ 19,500 $ 34,000 Cost of goods sold 585,100 632,500 Accounts receivable, net 37,400 57,400 Interest expense 7,900 13,000 Current notes receivable (trade) 9,100 7,200 Income tax expense 14,800 24,300 Merchandise inventory 84,440 132,500 Net income 162,200 210,400 Prepaid expenses 5,000 6,950 Basic earnings per share 4.51 5.11 Plant assets, net 290,000 304,400 Cash dividends per share 3.81 3.93 Total assets $ 445,440 $ 542,450 Beginning-of-year balance sheet data Liabilities and Equity Accounts receivable, net $ 29,800 $ 54,200 Current liabilities $ 61,340 $ 93,300 Current notes receivable (trade) 0 0 Long-term notes payable 80,800 101,000 Merchandise inventory 55,600 107,400 Common stock, $5 par value 180,000 206,000 Total assets 398,000 382,500 Retained earnings 123,300 142,150 Common stock, $5 par value 180,000 206,000 Total liabilities and equity $ 445,440 $ 542,450 Retained earnings 98,300 93,600 ________________________________________ Required: 1a. For both companies compute the (a) current ratio, (b) acid-test ratio, (c) accounts (including notes) receivable turnover, (d) inventory turnover, (e) days' sales in inventory, and (f) days' sales uncollected. (Do not round intermediate calculations.) 1b. Identify the company you consider to be the better short-term credit risk. For both companies compute the days' sales uncollected. (f) Days’ Sales Uncollected Company Choose Numerator: / Choose Denominator: X Days = Days’ Sales Uncollected Current receivables, net / Net Sales X 365 = days Barco / $ 770,000 X 365 = days Kyan / $ 880,200 X 365 = days
Answer #1
Barco Company Kyan Company
a)
Current Ratio
Current Assets/Current Liabilites
Current Assets
Cash $             19,500.00 34000
Accounts Receivable $             37,400.00 57400
Current Notes receivable(trade) $               9,100.00 7200
Inventory $             84,440.00 132500
Prepaid Expenses $               5,000.00 6950
Total Current Assets=(A) $         1,55,440.00 238050
Total Current Liabilites=(B) $             61,340.00 93300
Current Ratio=(A)/(B) 2.53:1 2.55:1
b)
Acid Test Ratio
Liquid Assets/Current Liabilites
Liquid Assets
Cash $             19,500.00 34000
Accounts Receivable $             37,400.00 57400
Current Notes receivable(trade) $               9,100.00 7200
Total Liquid Assets=(A) $             66,000.00 98600
Total Current Liabilites=(B) $             61,340.00 93300
Acid Test Ratio=(A)/(B) 1.08:1 1.06:1
c )
Accounts(inculding notes) receivable turnover
Sales=(A) $         7,70,000.00 $                   8,80,200.00
Beginning Accounts Receivable $             29,800.00 $                       54,200.00
Ending Accounts Receivable(Including Note) $             46,500.00 $                       64,600.00
Average Accounts Receivable=(Beginning Accounts Receivable+Ending Accounts Receivable/2)=($29800+$46500)/2,($54200+$64600)/2=(B) $             38,150.00 $                       59,400.00
Accounts Receivable Turnover=Sales/Average accounts receivable=(A)/(B) 20.18 14.82 Times
d)
Inventory Turnover=Cost of goods sold/Average Inventory
Cost of good sold=(A) $         5,85,100.00 $                   6,32,500.00
Beginning Inventory $             55,600.00 $                   1,07,400.00
Ending Inventory $             84,440.00 $                   1,32,500.00
Average Inventory=(Beginning Inventory+Ending Inventory)/2,($55600+$84440)/2,($107400+$132500)/2=(B) $             70,020.00 $                   1,19,950.00
Inventory Turnover=(A)/(B) 8.36 5.27 Times
e)
Days sales in Inventory=(365/Inventory Turnover)
(365/8.36),(365/5.27) 43.66 69.26 Days
f)
Days sales uncollected=(365/Accounts receivable turnover)
(365/20.18),(365/14.82) 18.08 24.63 Days